OpenStudy (anonymous):

lim e^-x +5/4^1/x as x approaches positive infinity

OpenStudy (zzr0ck3r):

\(\lim_{x\rightarrow \infty}e^{-x}+(\frac{5}{4})^{\frac{1}{x}}\)?

OpenStudy (zzr0ck3r):

@vicky38 ??

OpenStudy (anonymous):

no its plus 5 divide by 4 to the 1over x power

OpenStudy (zzr0ck3r):

\(\frac{e^{-x}+5}{4^{\frac{1}{x}}}\)

OpenStudy (anonymous):

lim e-x + 5 (l) lim 4 – x x® ∞ 41/x yes

OpenStudy (zzr0ck3r):

well \(e^{-x}=\frac{1}{e^x}\rightarrow \frac{1}{\infty}=0 \) as \(x\rightarrow \infty\) and \(4^{\frac{1}{x}}\rightarrow 4^{\frac{1}{\infty}}\rightarrow 4^0=1 \) as \(x\rightarrow \infty\) so we have \(\frac{0+5}{1}=5\)

OpenStudy (anonymous):

ignore the first mail the attached doc i am doing the letter k on it

OpenStudy (anonymous):

how would letter O work out

OpenStudy (zzr0ck3r):

\(\frac{2}{\infty} \rightarrow0\) so \(8\)

OpenStudy (anonymous):

thanks for the other problem had the same idea just did not know how to finish the thought

OpenStudy (anonymous):

what about letter L